Reports 1-1 of 1 Clear search Modify search
Detector Operation (Operations Report)
berni - 15:56 Monday 04 December 2023 (62631) Print this report
Operator Report - Morning shift

This morning I found the ITF locked in LN3 thus at 6:56 UTC I launched the daily calibration; ITF unlcoked during the calibration.

It relocked at 9:31 UTC and it remained locked until the end of the shift; the planned activity on ISc was carried out without issues.



At 14:08 we performed a step in PR CHROCC: 2023-12-04 14h08m34 UTC    berni 'OUT:DAC_V [CHROCC_V_DAC=2.35,ramp time=1800]' sent to TCS_CHROCC_PR

Images attached to this report
Comments to this report:
rolland - 18:13 Friday 08 December 2023 (62682) Print this report

Some results of "daily calibration" from December 1st, 4th and 8th

Injections to check hrec (estimation of bias)

The first four plots show, for a given actuator to inject hinj, the ratio hrec/hinj for different datasets, which are rather stable over the 3 different days.

Then, these dataset have been averaged (even if not everywhere compatible with statistics, mainly for f<30 Hz): see three examples in figures 5, 6 (for MirNE) and 7 (for PcalNE).

The 8th figure shows the superposition of the averaged results for the different actuators: they all see the same bias of h(t), with currently ~2% maximum differences in absolute amplitude.


Measurement of optical responses

The model for DARM sensing has been updated, and used for the analysis of some optical response measurements.

The model fitted on these data is based on the model proposed by Michal in logbook entry 62250.

A preliminary results are shown in figure 9 for optical response to a motion of NE mirror. Ann optical spring has been fixed with frequency of 4 Hz and Q = 10 (which is not exact looking the low frequency). The fitted parameters, for data from December 1st, are:

  • optical gain 1.6e9 W/m
  • frequency of the zero: 368.7 Hz
  • SR detuning angle: ~0.                    which gives a DCP at 368.7 Hz with Q = 0.50
  • extra-delay of -13 µs       -> this being quite close to 0 gives confidence in the data modeling and fit result.

The optical response when moving the WE mirror few minutes apparts is shown in figure 10, with the following parameters:

  • optical gain 1.7e9 W/m
  • frequency of the zero: 375.0 Hz
  • SR detuning angle: ~-7 deg.                    which gives a DCP at 374.3 Hz with Q = 0.502
  • extra-delay of -18 µs

More data from 4th and 8th December are still to be analysed to estimate if this fit is stable and can highlight optical response different from a simple pole (i.e. SR detuning).


NOTE: in the last two "daily calibration" (4 and 8 December), the ITF unlocked during the injections for the optical measurements (injections on DARM and on the different mirror actuators), but not always at the same injection.

Images attached to this comment
Search Help